THE ROTATIONAL SPECTRA OF IO X_{1} ^{2}Π_{3/2}, v ≤ 13 and X_{2} ^{2}Π_{1/2}, v ≤ 9

\underline{CHARLES E. MILLER}, Department of Chemistry, Haverford College, Haverford, PA 19041-1392; and EDWARD A. COHEN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099.

The rotational spectra of IO in vibrational states up to v = 13 in the X_{1} ^{2}Π_{3/2} state and up to v = 9 in the X_{2} ^{2}Π_{1/2} state have been observed in an O_{2} discharge over molecular I_{2}. In addition, I^{18}O has been observed for both the X_{1} and X_{2} states up to v = 5. All data have been analyzed simultaneously with fixed isotopic ratios among the constants. This extends the data set for the X_{1} state described last year at this meeting and provides the first high resolution data for the X_{2} state and for I^{18}O. An extensive set of parameters has been derived. These will be interpreted in terms of the electronic structure and the interatomic potential.